# Transaction Fee Market Design for Parallel Execution

### v1

@@ 4.2. Mechanisms without Easy Gas Estimation
We now switch gears towards mechanisms without Easy Gas Estimation that instead aim to achieve Scheduling Sensitivity and/or Efficiency. Intuitively, for a set of transactions $T$, such mechanisms should start from $v(T)$ and take each transaction’s marginal contribution towards $v(T)$ as its gas consumption. This, however, has to be done carefully, as, e.g., simply looking for each transaction $\textit{tx}\in T$ at $v(T)-v(T\setminus\{\textit{tx}\})$ does not help. Note that one can easily construct cases where removing any one transaction from $T$ does not change $v(T)$, so all reported numbers will be 0. However, all these numbers being zero does not mean that no transaction contributes to $v(T)$; it just means that we also have to consider removing multiple transactions to see the differences. This gives us the idea to look at the marginal contribution of each transaction $\textit{tx}\in T$ when added to each possible subset $S\subseteq T\setminus\{\textit{tx}\}$, and not just to $S=T\setminus\{\textit{tx}\}$ as before. Formally, we would like to compute tx’s gas consumption as an aggregate of $\left(v(S\cup\{\textit{tx}\})-v(S)\right)_{S\subseteq T\setminus\{\textit{tx}\}}$. We next present two mechanisms based on this idea: the Shapley (Definition 4.2) and Banzhaf (Definition 4.4) GCMs, inspired by corresponding concepts in cooperative game theory.
@@
(2) $\textsf{gas}_{T}^{S}(\textit{tx}):=$ $\frac{1}{n!}\sum_{\sigma}\left[v(P_{\textit{tx}}^{\sigma}\cup\{\textit{tx}\})-v(P_{\textit{tx}}^{\sigma})\right]$
(3) $=$ $\sum_{S\subseteq T\setminus\{\textit{tx}\}}\frac{|S|!\cdot(n-|S|-1)!}{n!}\left[v(S\cup\{\textit{tx}\})-v(S)\right].$
Here, $\sigma$ ranges over the $n!$ possible ways to order the transactions in $T$ and $P_{\textit{tx}}^{\sigma}$ denotes the set of transactions that precede tx in the order $\sigma$. The equality between Eqs. 2 and 3 follows by counting the number of orders $\sigma$ such that $P_{\textit{tx}}^{\sigma}=S$, which there are $|S|!\cdot(n-|S|-1)!$ of.
@@
The Shapley GCM satisfies Efficiency (Property 7).
@@
The Banzhaf GCM does not satisfy Efficiency (Property 7).
@@
5. Analysis of Our GCMs
In this section, we provide a detailed analysis of the normative properties of our GCMs. Our results are summarized in Table 1.
@@ 5. Analysis of Our GCMs
Resource Sensitivity (Property 1) $=$ $<$ $\leq$ $\leq$ $\leq$ $\leq$ $\leq$
Time Sensitivity (Property 2) $<$ $<$ $<$ $<$ $<$ $\leq$ $\leq$
Resource-Time Sensitivity (Property 3) $\leq$ $<$ $\leq$ $\leq$ $\leq$ $\leq$ $\leq$
@@
Transaction Bundling (Property 5) $=$ $\leq$ $\leq$ $\leq$ $<$
Scheduling Sensitivity (Property 6) $<$ $<$
Efficiency (Property 7)
Easy Gas Estimation (Property 8)
Poly-time Computable (Property 9) $S(v)$ $B(v)$ $v$ $v$ $v$
[…]
5.1. Mechanisms with Easy Gas Estimation
In this section, we analyze the current and Weighted Area GCMs. By definition, both satisfy Easy Gas Estimation (Property 8) and are Poly-time Computable (Property 9). Since they satisfy Easy Gas Estimation but are not constant GCMs, Theorem 3.4 implies that neither satisfies Scheduling Sensitivity (Property 6) nor Efficiency (Property 7). Furthermore, both mechanisms satisfy strict Time Sensitivity (Property 2): increasing the time $t$ of a transaction strictly increases its gas consumption. Similarly, both satisfy strict Set Inclusion (Property 4): since transactions’ gas consumptions are strictly positive, a strict superset of a given set of transactions consumes strictly more gas. For the remaining three properties, the two mechanisms behave slightly differently:
Resource Sensitivity (Property 1). The current mechanism ignores the set of resources $R$ that a transaction accesses, so replacing $R$ with a strict superset of it does not change the transaction’s gas consumption. Therefore, the current mechanism satisfies Resource Sensitivity with equality. On the other hand, the Weighted Area mechanism adds an extra term of $t\cdot w_{r}>0$ to the gas consumption for each additional resource $r$ added to $R$, so it satisfies strict Resource Sensitivity.
Resource-Time Sensitivity (Property 3). From the above results on whether our two mechanisms satisfy Resource Sensitivity and Time Sensitivity, Lemmas 3.1 and 3.2 allow us to conclude that the current mechanism satisfies Resource-Time Sensitivity, while the Weighted Area mechanism satisfies it strictly.
Transaction Bundling (Property 5). Concatenating two transactions with times $t_{1}$ and $t_{2}$ results in a transaction with time $t_{1}+t_{2}$. Since the current mechanism equates gas consumption with time, the bundled transaction has the same gas consumption as the two individual transactions combined. This implies that the current mechanism satisfies Transaction Bundling with equality. Finally, the Weighted Area mechanism satisfies Transaction Bundling, as shown below. If we restrict ourselves to bundling transactions with different resource sets, the property is strictly satisfied.
@@ 5.1. Mechanisms with Easy Gas Estimation
The Weighted Area GCM satisfies Transaction Bundling (Property 5). If we only consider bundling transactions with different resource sets, the property is strictly satisfied.
Proof.
@@
Consider two transactions $\textit{tx}_{1}\simeq(t_{1},R_{1})$ and $\textit{tx}_{2}\simeq(t_{2},R_{2})$. Let $\textit{tx}_{3}\simeq(t_{1}+t_{2},R_{1}\cup R_{2})$ be a transaction consisting of the concatenation of $\textit{tx}_{1}$ and $\textit{tx}_{2}$. We want to show that $\textsf{gas}^{\textit{WA}}(\textit{tx}_{1})+\textsf{gas}^{\textit{WA}}(\textit{tx}_{2})\leq\textsf{gas}^{\textit{WA}}(\textit{tx}_{3})$. By definition, this amounts to:
(4) $t_{1}\cdot\left(1+\sum_{r\in R_{1}}w_{r}\right)+t_{2}\cdot\left(1+\sum_{r\in R_{2}}w_{r}\right)\leq(t_{1}+t_{2})\cdot\left(1+\sum_{r\in R_{1}\cup R_{2}}w_{r}\right)$
@@
Which is true because $\sum_{r\in R_{i}}w_{r}\leq\sum_{r\in R_{1}\cup R_{2}}w_{r}$ for $i\in\{1,2\}$. Because the weights are strictly positive, equality occurs if and only if $R_{1}=R_{1}\cup R_{2}$ and $R_{2}=R_{1}\cup R_{2}$; i.e., $R_{1}=R_{2}$. Hence, the property is satisfied strictly if we restrict bundling transactions to cases where $R_{1}\neq R_{2}$. ∎
5.2. Mechanisms without Easy Gas Estimation
@@
In this section, we analyze the Shapley, Banzhaf, TPM, ESM, and XSM GCMs. By definition, all of them require knowledge of $T$ to compute $\textsf{gas}_{T}(\textit{tx})$, so they do not satisfy Easy Gas Estimation (Property 8). Next, we examine each of the remaining properties individually and analyze whether they hold for our five mechanisms.
@@
Poly-time Computable (Property 9). The TPM, ESM, and XSM GCMs are all Poly-time Computable whenever determining the makespan of a given set of transactions under the chosen scheduler is feasible in polynomial time, i.e., when $v(T)$ can be computed in polynomial time. [^15] This is essentially never true for non-trivial makespan minimization problems. In particular, if we restrict our attention to the case of unit-length transactions (i.e., with $t=1$) and infinitely many threads $n=\infty$, then checking for the existence of a schedule with makespan $c$ corresponds to checking whether the intersection graph of the transactions (i.e., where edges correspond to transactions with intersecting resource sets) is $c$-colorable. For $c=3$, this is well-known to be NP-complete, but this result is for general graphs. However, there is a straightforward way to model any graph $G=(V,E)$ as an intersection graph of transactions: vertices $v\in V$ are transactions and for every edge $(u,v)\in E$, create a new resource $r$ and add it to the resource sets of transactions $u$ and $v$. In fact, the problems are all equally difficult to computing such makespans. In contrast, computing gas consumptions under the Shapley and Banzhaf GCMs hits another hurdle: it is no longer enough to be able to efficiently compute $v(T)$, but instead, one needs to be able to compute the Shapley or Banzhaf values of the transactions, involving aggregating over $(v(T^{\prime}))_{T^{\prime}\subseteq T}$. Hence, computing Shapley and Banzhaf values is, in general, NP-hard and #P-complete (Deng and Papadimitriou, 1994; Faliszewski and Hemaspaandra, 2009; Matsui and Matsui, 2000, 2001; Prasad and Kelly, 1990). Moreover, no deterministic polynomial-time algorithm can approximate the Shapley values within a constant factor unless P = NP. Using randomization could, in principle, circumvent this: to approximate an average consisting of exponentially many terms, sample polynomially many uniformly at random, and take their average. However, randomization would not be appropriate for the blockchain context, and this still runs into the problem of computing the sampled terms, which is as hard as evaluating $v$ a constant number of times. See (Chalkiadakis et al., 2012) for an ampler discussion of computing Shapley and Banzhaf values. Note that the previous results mostly pertain to functions $v$ of a different shape than ours (i.e., not related to scheduling). We have not attempted to show that hardness is retained in our context, but expect this to be true. Note still that because the Shapley GCM is Efficient, it can be used to compute $v(T)$ by adding up the gas consumptions of all transactions in $T$, meaning that the Shapley GCM is at least as difficult to compute as $v(T)$. However, this simple reduction no longer works for the Banzhaf GCM.
Efficiency (Property 7). We have already seen that the Shapley GCM is Efficient (Lemma 4.3) while the Banzhaf GCM is not (Lemma 4.5). Moreover, by adding up the gas consumption, one can immediately see from the definitions that TPM and ESM are Efficient, while XSM is not.
@@
Scheduling Sensitivity (Property 6). The ESM and XSM GCMs can be easily seen to strictly satisfy Scheduling Sensitivity. This is because, under both mechanisms, given a set of transactions $T$, the gas consumption of any transaction in $T$ is computed as $f(|T|)\cdot v(T)$, where $f$ is either $\frac{1}{x}$ or $\frac{1}{3^{x}}$. Notably, the first factor depends only on $|T|$, so if a transaction in $T$ were modified in a way that increases the makespan, this increase would also be reflected proportionally in its gas consumption. In contrast, the Shapley, Banzhaf, and TPM GCMs do not satisfy Scheduling Sensitivity (proven in Lemma 5.2 below). This may be particularly surprising for the Shapley and Banzhaf GCMs, as they were specifically designed to account for marginal increases in makespan. The catch is that Scheduling Sensitivity considers replacing a transaction tx with another one leading to an increase in makespan. However, some elements in $\left(v(S\cup\{\textit{tx}\})-v(S)\right)_{S\subseteq T\setminus\{\textit{tx}\}}$ may still decrease as a result (except for $S=T\setminus\{\textit{tx}\}$, for which we assumed an increase). Because both the Shapley and Banzhaf values of tx take a weighted average over these values, the average might still decrease, which is what happens in the example in the lemma below.
@@
The Shapley, Banzhaf and TPM GCMs do not satisfy Scheduling Sensitivity (Property 6).
Proof.
Consider the set of transactions $T=\{\textit{tx}_{1},\textit{tx}_{2}\}$ and two additional transactions $\textit{tx}_{3}$ and $\textit{tx}_{4}$ such that $\textit{tx}_{1}\simeq(1,\{r_{1}\})$, $\textit{tx}_{2}\simeq(3,\{r_{2}\})$, $\textit{tx}_{3}\simeq(2,\{r_{1}\})$, and $\textit{tx}_{4}\simeq(1,\{r_{2}\})$. Then, for any number of threads $n\geq 2$ and the optimal scheduler we have $v(T\cup\{\textit{tx}_{3}\})=3<4=v(T\cup\{\textit{tx}_{4}\})$. However:
@@
So, the Shapley, Banzhaf, and TPM GCMs all violate Scheduling Sensitivity in this case. ∎
Transaction Bundling (Property 5). We find that the Banzhaf, TPM and XSM GCMs satisfy Transaction Bundling, only XSM satisfying it strictly, while the Shapley and ESM GCMs do not satisfy the property. We prove these facts in the following 5 lemmas.
@@
The Shapley GCM does not satisfy Transaction Bundling (Property 5).
@@
The Banzhaf GCM satisfies Transaction Bundling (Property 5).
@@
The TPM GCM satisfies Transaction Bundling (Property 5).
@@
The ESM GCM does not satisfy Transaction Bundling (Property 5).
@@
The XSM GCM strictly satisfies Transaction Bundling (Property 5).
Set Inclusion (Property 4). We find that the TPM and ESM GCMs satisfy Set Inclusion, while the Shapley, Banzhaf and XSM GCMs do not satisfy the property. We prove these facts in the following 5 lemmas.
@@
The Shapley GCM does not satisfy Set Inclusion (Property 4).
@@
The Banzhaf GCM does not satisfy Set Inclusion (Property 4).
@@
The TPM GCM satisfies Set Inclusion (Property 4).
@@
The ESM GCM satisfies Set Inclusion (Property 4).
@@
The XSM GCM does not satisfy Set Inclusion (Property 4).
Resource-Time Sensitivity (Property 3). All five mechanisms satisfy this property. For the ESM and XSM GCMs, this is an immediate consequence of property (S3). For the Shapley and Banzhaf GCMs, one can see this by recalling that they compute the gas consumption of a transaction $\textit{tx}\in T$ as a weighted average over $\left(v(S\cup\{\textit{tx}\})-v(S)\right)_{S\subseteq T\setminus\{\textit{tx}\}}$. Hence, by property (S3), when tx is replaced with some $^{\prime}\gtrsim\textit{tx}$ tx, no term in the previous decreases, so their weighted average also does not decrease. Finally, for the TPM GCM, we show this in the lemma after the next paragraph.
Resource Sensitivity (Property 1) and Time Sensitivity (Property 2). Because all five mechanisms satisfy Resource-Time Sensitivity, by Lemma 3.1, they also all satisfy Resource Sensitivity and Time Sensitivity. Out of the five mechanisms, the Shapley, Banzhaf, and TPM GCMs satisfy the property strictly. For the first two, this is because when $t(\textit{tx})$ increases, at least one term in $\left(v(S\cup\{\textit{tx}\})-v(S)\right)_{S\subseteq T\setminus\{\textit{tx}\}}$ strictly increases, namely the term for $S=\varnothing$, while no terms decrease by property (S3). Last, for the TPM GCM, we show this in the lemma below.
@@
The TPM GCM satisfies Resource-Time Sensitivity (Property 3) and strict Time Sensitivity (Property 2).
6. Towards a Fee Market for Parallel Execution

### v2

@@ 4.2 Mechanisms without Easy Gas Estimation
We now switch gears towards mechanisms without Easy Gas Estimation that instead aim to achieve Scheduling Monotonicity and/or Efficiency. Intuitively, for a set of transactions $T$, such mechanisms should start from $v(T)$ and take each transaction’s marginal contribution towards $v(T)$ as its gas consumption. This, however, has to be done carefully, as, e.g., simply looking for each transaction $\textit{tx}\in T$ at $v(T)-v(T\setminus\{\textit{tx}\})$ does not help. Note that one can easily construct cases where removing any one transaction from $T$ does not change $v(T)$, so all reported numbers will be 0. However, all these numbers being zero does not mean that no transaction contributes to $v(T)$; it just means that we also have to consider removing multiple transactions to see the differences. This gives us the idea to look at the marginal contribution of each transaction $\textit{tx}\in T$ when added to each possible subset $S\subseteq T\setminus\{\textit{tx}\}$, and not just to $S=T\setminus\{\textit{tx}\}$ as before. Formally, we would like to compute tx’s gas consumption as an aggregate of $\left(v(S\cup\{\textit{tx}\})-v(S)\right)_{S\subseteq T\setminus\{\textit{tx}\}}$. We next present two mechanisms based on this idea: the Shapley (Definition˜4.2) and Banzhaf (Definition˜4.4) GCMs, inspired by corresponding concepts in cooperative game theory.
@@
$\textsf{gas}_{T}^{S}(\textit{tx}):=$ $\frac{1}{n!}\sum_{\sigma}\left[v(P_{\textit{tx}}^{\sigma}\cup\{\textit{tx}\})-v(P_{\textit{tx}}^{\sigma})\right]$ (2)
$=$ $\sum_{S\subseteq T\setminus\{\textit{tx}\}}\frac{|S|!\cdot(n-|S|-1)!}{n!}\left[v(S\cup\{\textit{tx}\})-v(S)\right].$ (3)
Here, $\sigma$ ranges over the $n!$ possible ways to order the transactions in $T$ and $P_{\textit{tx}}^{\sigma}$ denotes the set of transactions that precede tx in the order $\sigma$. The equality between Equations˜2 and 3 follows by counting the number of orders $\sigma$ such that $P_{\textit{tx}}^{\sigma}=S$, which there are $|S|!\cdot(n-|S|-1)!$ of.
@@
The Shapley GCM satisfies Efficiency (Property˜7).
@@
The Banzhaf GCM does not satisfy Efficiency (Property˜7).
@@
5 Analysis of Our GCMs
In this section, we provide a detailed analysis of the normative properties of our GCMs. Our results are summarized in Table˜1.
@@ 5 Analysis of Our GCMs
Storage Key Monotonicity (Property˜1) == << $\leq$ $\leq$ $\leq$ $\leq$ $\leq$
Time Monotonicity (Property˜2) << << << << << $\leq$ $\leq$
Storage Key-Time Monotonicity (Property˜3) $\leq$ << $\leq$ $\leq$ $\leq$ $\leq$ $\leq$
@@
Transaction Bundling (Property˜5) == $\leq$ $\leq$ $\leq$ <<
Scheduling Monotonicity (Property˜6) << <<
Efficiency (Property˜7)
Easy Gas Estimation (Property˜8)
Poly-time Computable (Property˜9) $S(v)$ $B(v)$ $v$ $v$ $v$
Table 1: Comparison of GCMs based on their adherence to the defined properties. <\,<\, indicates that the mechanism strictly satisfies the property, =\,=\, indicates trivial satisfaction (by equality), and $\leq\,$ indicates satisfaction (not necessarily strict). A $\checkmark$” means the property is satisfied; $\times$” means it is not satisfied. For computational complexity, “vv” means the mechanism is as hard to compute as $v(\cdot)$ itself, “S(v)S(v)” means it is as hard as computing Shapley values for $v$, and “B(v)B(v)” means it is as hard as computing Banzhaf values for $v$.
5.1 Mechanisms with Easy Gas Estimation
In this section, we analyze the current and Weighted Area GCMs. By definition, both satisfy Easy Gas Estimation (Property˜8) and are Poly-time Computable (Property˜9). Since they satisfy Easy Gas Estimation but are not constant GCMs, Theorem˜3.4 implies that neither satisfies Scheduling Monotonicity (Property˜6) nor Efficiency (Property˜7). Furthermore, both mechanisms satisfy strict Time Monotonicity (Property˜2): increasing the time $t$ of a transaction strictly increases its gas consumption. Similarly, both satisfy strict Set Inclusion (Property˜4): since transactions’ gas consumptions are strictly positive, a strict superset of a given set of transactions consumes strictly more gas. For the remaining three properties, the two mechanisms behave slightly differently:
Storage Key Monotonicity (Property˜1). The current mechanism ignores the set of storage keys $K$ that a transaction accesses, so replacing $K$ with a strict superset of it does not change the transaction’s gas consumption. Therefore, the current mechanism satisfies Storage Key Monotonicity with equality. On the other hand, the Weighted Area mechanism adds an extra term of $t\cdot w_{k}>0$ to the gas consumption for each additional storage key $k$ added to $K$, so it satisfies strict Storage Key Monotonicity.
Storage Key-Time Monotonicity (Property˜3). From the above results on whether our two mechanisms satisfy Storage Key Monotonicity and Time Monotonicity, Lemmas˜3.1 and 3.2 allow us to conclude that the current mechanism satisfies Storage Key-Time Monotonicity, while the Weighted Area mechanism satisfies it strictly.
Transaction Bundling (Property˜5). Concatenating two transactions with times $t_{1}$ and $t_{2}$ results in a transaction with time $t_{1}+t_{2}$. Since the current mechanism equates gas consumption with time, the bundled transaction has the same gas consumption as the two individual transactions combined. This implies that the current mechanism satisfies Transaction Bundling with equality. Finally, the Weighted Area mechanism satisfies Transaction Bundling, as shown below. If we restrict ourselves to bundling transactions with different storage key sets, the property is strictly satisfied.
@@ 5.1 Mechanisms with Easy Gas Estimation
The Weighted Area GCM satisfies Transaction Bundling (Property˜5). If we only consider bundling transactions with different storage key sets, the property is strictly satisfied.
Proof 5.2.
@@
Consider two transactions $\textit{tx}_{1}\simeq(t_{1},K_{1})$ and $\textit{tx}_{2}\simeq(t_{2},K_{2})$. Let $\textit{tx}_{3}\simeq(t_{1}+t_{2},K_{1}\cup K_{2})$ be a transaction consisting of the concatenation of $\textit{tx}_{1}$ and $\textit{tx}_{2}$. We want to show that $\textsf{gas}^{\textit{WA}}(\textit{tx}_{1})+\textsf{gas}^{\textit{WA}}(\textit{tx}_{2})\leq\textsf{gas}^{\textit{WA}}(\textit{tx}_{3})$. By definition, this amounts to:
$t_{1}\cdot\left(1+\sum_{k\in K_{1}}w_{k}\right)+t_{2}\cdot\left(1+\sum_{k\in K_{2}}w_{k}\right)\leq(t_{1}+t_{2})\cdot\left(1+\sum_{k\in K_{1}\cup K_{2}}w_{k}\right)$ (4)
@@
Which is true because $\sum_{k\in K_{i}}w_{k}\leq\sum_{k\in K_{1}\cup K_{2}}w_{k}$ for $i\in\{1,2\}$. Because the weights are strictly positive, equality occurs if and only if $K_{1}=K_{1}\cup K_{2}$ and $K_{2}=K_{1}\cup K_{2}$; i.e., $K_{1}=K_{2}$. Hence, the property is satisfied strictly if we restrict bundling transactions to cases where $K_{1}\neq K_{2}$.
5.2 Mechanisms without Easy Gas Estimation
@@
In this section, we analyze the Shapley, Banzhaf, TPM, ESM, and XSM GCMs. By definition, all of them require knowledge of $T$ to compute $\textsf{gas}_{T}(\textit{tx})$, so they do not satisfy Easy Gas Estimation (Property˜8). Next, we examine each of the remaining properties individually and analyze whether they hold for our five mechanisms.
@@
Poly-time Computable (Property˜9). The TPM, ESM, and XSM GCMs are all Poly-time Computable whenever determining the makespan of a given set of transactions under the chosen scheduler is feasible in polynomial time, i.e., when $v(T)$ can be computed in polynomial time. [^17] Notably, this is essentially never true for non-trivial makespan minimization problems. In particular, if we restrict our attention to the case of unit-length transactions (i.e., with $t=1$) and infinitely many threads $n=\infty$, then checking for the existence of a schedule with makespan $c$ corresponds to checking whether the intersection graph of the transactions (i.e., where edges correspond to transactions with intersecting storage key sets) is $c$-colorable. For $c=3$, this is well-known to be NP-complete, but this result is for general graphs. However, there is a straightforward way to model any graph $G=(V,E)$ as an intersection graph of transactions: vertices $v\in V$ are transactions and for every edge $(u,v)\in E$, create a new storage key $k$ and add it to the storage key sets of transactions $u$ and $v$. In fact, the problems are all equally difficult to computing such makespans. In contrast, computing gas consumptions under the Shapley and Banzhaf GCMs hits another hurdle: it is no longer enough to be able to efficiently compute $v(T)$, but instead, one needs to be able to compute the Shapley or Banzhaf values of the transactions, involving aggregating over $(v(T^{\prime}))_{T^{\prime}\subseteq T}$. Hence, computing Shapley and Banzhaf values is, in general, NP-hard and #P-complete [18, 23, 40, 41, 51]. Moreover, no deterministic polynomial-time algorithm can approximate the Shapley values within a constant factor unless P = NP. Using randomization could, in principle, circumvent this: to approximate an average consisting of exponentially many terms, sample polynomially many uniformly at random, and take their average. One could imagine implementing this in a blockchain via a VRF or similar mechanism to ensure unbiased randomness. Still, this comes with a notable downside: the accuracy of the computed gas consumption may be hard to predict in advance and might vary wildly. Moreover, even with randomization, one could still run into issues computing the sampled terms, which is as hard as evaluating $v$ a constant number of times. See [14] for an ampler discussion of computing Shapley and Banzhaf values. Note that the previous results mostly pertain to functions $v$ of a different shape than ours (i.e., not related to scheduling). We have not attempted to show that hardness is retained in our context, but expect this to be true. Note still that because the Shapley GCM is Efficient, it can be used to compute $v(T)$ by adding up the gas consumptions of all transactions in $T$, meaning that the Shapley GCM is at least as difficult to compute as $v(T)$. However, this simple reduction no longer works for the Banzhaf GCM.
Efficiency (Property˜7). We have already seen that the Shapley GCM is Efficient (Lemma˜4.3) while the Banzhaf GCM is not (Lemma˜4.5). Moreover, by adding up the gas consumption, one can immediately see from the definitions that TPM and ESM are Efficient, while XSM is not. We take this occasion to also note that any GCM can be made Efficient by appropriately scaling its output. Most notably, applying this to the Banzhaf GCM yields a mechanism based on the well-known normalized Banzhaf values, though the resulting mechanism otherwise seems rather poorly behaved.
@@
Scheduling Monotonicity (Property˜6). The ESM and XSM GCMs can be easily seen to strictly satisfy Scheduling Monotonicity. This is because, under both mechanisms, given a set of transactions $T$, the gas consumption of any transaction in $T$ is computed as $f(|T|)\cdot v(T)$, where $f$ is either $\frac{1}{x}$ or $\frac{1}{3^{x}}$. Notably, the first factor depends only on $|T|$, so if a transaction in $T$ were modified in a way that increases the makespan, this increase would also be reflected proportionally in its gas consumption. In contrast, the Shapley, Banzhaf, and TPM GCMs do not satisfy Scheduling Monotonicity (proven in Lemma˜5.3 below). This may be particularly surprising for the Shapley and Banzhaf GCMs, as they were specifically designed to account for marginal increases in makespan. The catch is that Scheduling Monotonicity considers replacing a transaction tx with another one leading to an increase in makespan. However, some elements in $\left(v(S\cup\{\textit{tx}\})-v(S)\right)_{S\subseteq T\setminus\{\textit{tx}\}}$ may still decrease as a result (except for $S=T\setminus\{\textit{tx}\}$, for which we assumed an increase). Because both the Shapley and Banzhaf values of tx take a weighted average over these values, the average might still decrease, which is what happens in the example in the lemma below.
@@
The Shapley, Banzhaf and TPM GCMs do not satisfy Scheduling Monotonicity (Property˜6).
Proof 5.4.
Consider the set of transactions $T=\{\textit{tx}_{1},\textit{tx}_{2}\}$ and two additional transactions $\textit{tx}_{3}$ and $\textit{tx}_{4}$ such that $\textit{tx}_{1}\simeq(1,\{k_{1}\})$, $\textit{tx}_{2}\simeq(3,\{k_{2}\})$, $\textit{tx}_{3}\simeq(2,\{k_{1}\})$, and $\textit{tx}_{4}\simeq(1,\{k_{2}\})$. Then, for any number of threads $n\geq 2$ and the optimal scheduler we have $v(T\cup\{\textit{tx}_{3}\})=3<4=v(T\cup\{\textit{tx}_{4}\})$. However:
@@
So, the Shapley, Banzhaf, and TPM GCMs all violate Scheduling Monotonicity in this case.
Transaction Bundling (Property˜5). We find that the Banzhaf, TPM and XSM GCMs satisfy Transaction Bundling, only XSM satisfying it strictly, while the Shapley and ESM GCMs do not satisfy the property. We prove these facts in the following 5 lemmas.
@@
The Shapley GCM does not satisfy Transaction Bundling (Property˜5).
@@
The Banzhaf GCM satisfies Transaction Bundling (Property˜5).
@@
The TPM GCM satisfies Transaction Bundling (Property˜5).
@@
The ESM GCM does not satisfy Transaction Bundling (Property˜5).
@@
The XSM GCM strictly satisfies Transaction Bundling (Property˜5).
Set Inclusion (Property˜4). We find that the TPM and ESM GCMs satisfy Set Inclusion, while the Shapley, Banzhaf and XSM GCMs do not satisfy the property. We prove these facts in the following 5 lemmas.
@@
The Shapley GCM does not satisfy Set Inclusion (Property˜4).
@@
The Banzhaf GCM does not satisfy Set Inclusion (Property˜4).
@@
The TPM GCM satisfies Set Inclusion (Property˜4).
@@
The ESM GCM satisfies Set Inclusion (Property˜4).
@@
The XSM GCM does not satisfy Set Inclusion (Property˜4).
Storage Key-Time Monotonicity (Property˜3). All five mechanisms satisfy this property. For the ESM and XSM GCMs, this is an immediate consequence of property (S3). For the Shapley and Banzhaf GCMs, one can see this by recalling that they compute the gas consumption of a transaction $\textit{tx}\in T$ as a weighted average over $\left(v(S\cup\{\textit{tx}\})-v(S)\right)_{S\subseteq T\setminus\{\textit{tx}\}}$. Hence, by property (S3), when tx is replaced with some $\textit{tx}^{\prime}\gtrsim\textit{tx}$, no term in the previous decreases, so their weighted average also does not decrease. Finally, for the TPM GCM, we show this in the lemma after the next paragraph*.
Storage Key Monotonicity (Property˜1) and Time Monotonicity (Property˜2). Because all five mechanisms satisfy Storage Key-Time Monotonicity, by Lemma˜3.1, they also all satisfy Storage Key Monotonicity and Time Monotonicity. Out of the five mechanisms, the Shapley, Banzhaf, and TPM GCMs satisfy the property strictly. For the first two, this is because when $t(\textit{tx})$ increases, at least one term in $\left(v(S\cup\{\textit{tx}\})-v(S)\right)_{S\subseteq T\setminus\{\textit{tx}\}}$ strictly increases, namely the term for $S=\varnothing$, while no terms decrease by property (S3). Last, for the TPM GCM, we show this in the lemma below.
@@
The TPM GCM satisfies Storage Key-Time Monotonicity (Property˜3) and strict Time Monotonicity (Property˜2).
6 Towards a Fee Market for Parallel Execution